OpenStudy (anonymous):

If E and F are independent events with P(E)=.50 and P(F)=.10 , find:

OpenStudy (anonymous):

P(E I F) P(F I E) P(E n F) P(E u F)

OpenStudy (zarkon):

.5 .1 .05 .55

OpenStudy (zarkon):

\[P(E|F)=P(E)\] \[P(F|E)=P(F)\] \[P(E\cap F)=P(E)\cdot P(F)\] \[P(E\cup F)=P(E)+P(F)-P(E\cap F)\]